# A Quantum-Driven Evolutionary Framework for Solving High-Dimensional Sharpe Ratio Portfolio Optimization

### v1

@@ 3.2. Our proposed QHDE
3.2.3. dynamic elite pool with Cauchy-Gaussian mixed disturbance strategy
@@
In the later iterations of the DE, population individuals often tend to cluster around the current best solution due to rapid assimilation. If the current solution is not globally optimal, the population may prematurely fall into a local optimum, leading to a stagnation of the search. To address this issue, this paper proposes a dynamic elite pool and Cauchy-Gaussian mixed disturbance strategy. This strategy enhances the algorithm’s global search ability and diversity of solutions by constructing an elite pool (Elitep) consisting of the top-performing individuals, thereby avoiding the algorithm from being trapped in local optima. The specific procedure is as follows:
@@
First, based on the fitness evaluations of the population, the three highest fitness individuals, denoted as $PZ1$, $PZ2$, and $PZ3$, are selected. Then the arithmetic mean of their positions is calculated to obtain the average position $PZm$:
(31) $PZm=\frac{(PZ1+PZ2+PZ3)}{3}$
@@
Cauchy-Gaussian mutation is a hybrid mutation method that combines the long-tail characteristics of the Cauchy distribution with the concentration tendency of the Gaussian distribution. The purpose of this strategy is to apply strategic random perturbations to the global best individual, exploring potentially unknown beneficial regions around the global optimum, while maintaining the ”memory” of the current best solution. After combining the dynamic elite pool and Cauchy-Gaussian perturbation strategies, the update formula for the new individual’s position is as follows:
(32) $x^{t+1}_{\text{best}}=Elite\times[\rho\times\text{Cauchy}(0,1)+(1-\rho)\times\text{Gauss}(0,1)]$
where $x^{t+1}_{\text{best}}$ represents the global best individual in the $(t+1)$ generation. $\rho$ denotes a weight coefficient that balances the contributions of the Cauchy distributions and Gaussian distributions in the mutation operation, with values ranging from $[0,1]$. The choice of this parameter affects the randomness and distribution of the mutation. $\text{Cauchy}(0,1)$, with its long tails, introduces large perturbations that help escape local optima and enhance global search. In contrast, $\text{Gauss}(0,1)$ concentrates around the mean, enabling stable, fine-grained local search. The parameter $\rho$ balances the influence of these two distributions, allowing the mutation to adjust between global exploration and local exploitation.
@@
7:   Randomly select $x_{r_{1}},x_{r_{2}},x_{r_{3}}$ from the population and calculate $v_{i}$ using (13)
@@
9:   Generate $x_{new}$ using (30)
10:   if $X_{i}$ surpasses $u_{i}$ then
11:     $u_{i}=x_{new}$
12:   end if
13:   if $f(u_{i})\leq f(x_{i})$ then
14:     Insert $u_{i}$ into new generation $P_{new}$
15:   else
16:     Insert $x_{i}$ into new generation $P_{new}$
17:   end if
18:  end for
@@
19:  Select the fittest individuals $PZ1,PZ2,PZ3$ and calculate PZmPZm using (31) to form the elite pool
20:  Generate $x_{\text{best}}^{t+1}$ using (32)
21:  if $x_{\text{best}}^{t+1}$ surpasses $pos_{\text{best}}$ then
22:   $pos_{\text{best}}=x_{\text{best}}^{t+1}$
23:  end if
24:  Update best global position $pos_{\text{best}}$ and its fitness value $fit_{\text{best}}$
25:  $P=P_{new}$
26:  $t=t+1$
27:end while
28:return $pos_{\text{best}}$ and $fit_{\text{best}}$
3.3. Time and space complexity analysis
@@
The computational complexity of QHDE mainly depends on population initialization and core operations such as fitness evaluation and iterative updates. With population size $N$, iteration limit $T_{\max}$, and dimension $D$, the original complexity is $O(T_{\max}\texttimes N\texttimes D)$. The improved QHDE introduces good-point-set chaos-reverse initialization, quantum evolution with information exchange, and a dynamic elite pool with Cauchy–Gaussian disturbance. Initialization operations (good-point-set generation, chaotic mapping, and reverse learning) each cost $O(N\texttimes D)$ and are executed once. Quantum evolution updates and the elite-pool disturbance strategy both require $O(N\texttimes D)$ per iteration, giving $O(T_{\max}\texttimes N\texttimes D)$. Thus, the total complexity remains $O(T_{\max}\texttimes N\texttimes D)$, consistent with the original algorithm.
@@
The space complexity of DE depends on storing the population, fitness values, and candidate solutions. With population size $N$ and dimension $D$, storing positions and temporary candidates requires $O(N\texttimes D)$, while fitness values need $O(N)$. Thus, the overall complexity is $O(N\texttimes D)$. In the improved QHDE, storing good-point sets, reverse individuals, quantum states, and the elite pool also requires $O(N\texttimes D)$, so the total space complexity remains $O(N\texttimes D)$.
4. Experiments and comprehensive analysis
@@
This section provides a detailed discussion of the experimental results obtained by applying QHDE to numerical optimization tasks and portfolio selection problems. Firstly, we conduct numerical optimization tests using the CEC 2020 and CEC 2022 test suite to preliminarily evaluate the performance of the algorithm in optimizing multimodal functions and high-dimensional problems. Next, we apply QHDE to four portfolio selection problems (involving 20, 40, 60, and 80 stocks, respectively) and compare its performance with seven metaheuristic algorithms that are widely employed in this field to validate the superiority of QHDE in portfolio optimization. The simulations were conducted on a Windows 11 platform with a 64-bit operating system. The analyses were carried out using MATLAB 2023b on a system powered by an AMD Ryzen 7 4800H CPU at 2.30 GHz and 16 GB of RAM.
4.1. Test functions and comparison algorithms
@@
To initially assess the performance of the proposed QHDE, we utilized the CEC 2020 test suite (Liang et al., 2019) and CEC 2022 (Mohamed et al., 2020), which contain diverse, high-dimensional, constrained, and multimodal optimization problems closely mirroring real portfolio-optimization challenges. Their emphasis on non-convex objectives, multiple local optima, and complex constraints makes them strong proxies for balancing return–risk trade-offs. These benchmarks therefore provide a rigorous and representative basis for assessing QHDE’s effectiveness.
@@
The performance of QHDE was systematically compared against seven well-known algorithms. These are categorized into two distinct groups for a structured analysis:
State-of-the-art algorithm: Self-adaptive spherical search algorithm (Kumar et al., 2020) (SASS), L-SHADE for Constrained Optimization with Levy Flights (Gurrola-Ramos et al., 2020) (COLSHADE), Simple Modification in Covariance Matrix Adaptation Evolution Strategy (Ros and Hansen, 2008) (sCMAES), Hierarchical Self Evolutionary PSO (Wei et al., 2025) (HSEPSO), Niching Swarm Dynamic Optimization with TCMAES (Fujita et al., 2025) (NDSOT).
DE and its improved variants: This includes the original Differential Evolution algorithm (Das and Suganthan, 2010) (DE), the Advanced Differential Evolution (Abbasi et al., 2024) (ADE).
Table 1 provides an overview of the parameter settings for eight different MH algorithms.
Table 1. Parameter configurations for competing algorithms
@@
QHDE $\beta$, $P_{c}$, $w_{1}$, $w_{2}$, $c_{1}$, $c_{2}$ 0.5, 0.1, 0.5, 1, 1.5, 1.5
4.2. Quantitative evaluation
@@
This section provides a comprehensive evaluation of the QHDE’s performance using the CEC 2020 and CEC 2022 test suite. To ensure consistency, the experimental parameters were standardized with a population size of 30 and a maximum of 500 iterations, and each algorithm was executed 30 independent runs.
To better illustrate the performance of each algorithm, we summarize the ranking distribution of each algorithm, as shown in Figure  2. In the 10-dimensional benchmarks, QHDE achieved the best performance on two functions, ranked second on four, and third on two. Consequently, it consistently placed within the top three in 80% of the test instances, and didn’t exhibit the worst performance on any problem, demonstrating exceptional competitiveness and optimization precision. In contrast, although SASS and sCMAES secured the top rank in slightly more instances (one additional count), they exhibited significant performance volatility, with over half of their rankings classified as ’Other’ or lower. Notably, while the rankings of the original DE were predominantly concentrated below the third position, QHDE successfully elevated its standing to the first and second tiers. This substantial improvement substantiates that the incorporated quantum evolutionary strategy effectively enhances the algorithm’s capability to escape local optima and accelerates convergence. Overall, QHDE delivers not only high-quality solutions but also exceptional stability across different functions, proving to be the most robust algorithm in this comparative study.
@@
Figure  3 presents the Friedman rank radar plot of all participating algorithms on the CEC 2022 benchmark suite. As evident in Figure  3(a), QHDE (in red) consistently achieves lower rankings across the majority of test functions, forming a compact shape near the center of the plot. With the increase in dimensionality, the superiority of QHDE becomes even more pronounced. In Figure  3(b), its compactness is further accentuated as the curve closely hugs the central region, indicating that QHDE maintains top-tier performance across diverse problem types with minimal volatility. Notably, QHDE ranks first on functions F2, F3, F7, F8, F10, F11, and F12, which are known for their high complexity and multimodality. The overall rank of QHDE is the lowest among all contenders, further confirming its superior and robust optimization capability across a wide range of scenarios.
Figure 2. Ranking distribution of different algorithms on CEC 2020 (Dim=10).
Two bar charts showing ranking distributions for eight algorithms across different CEC 2020 functions with dimensions 10 and 20.
(a)                         (b) Radar chart comparing algorithm performance on 12 functions plus mean in CEC 2022 benchmark, dimension 20.
Figure 3. Ranking distribution of different algorithms on CEC 2022. (a) Dim = 10, (b) Dim = 20.
@@
As clearly shown in Figure  4, the boxplot analysis of various algorithms on representative functions reveals a significant stability advantage of QHDE. The box of the QHDE boxplot is relatively narrow, indicating a highly concentrated data distribution. This suggests that, across multiple experimental runs, the results generated by QHDE exhibit minimal dispersion, and the search process remains stable. For instance, when dealing with different types of CEC 2020 functions, regardless of the complexity and dimensionality of the functions, QHDE can precisely adjust its search direction and strategy by virtue of its the three embedded improvement mechanisms, avoid getting trapped in local optima, thereby continuously and stably output high-quality solutions.In contrast, for algorithms such as SASS, COLSHADE, HSEPSO and NDSOT, the overall positions of the boxplots are relatively higher. This suggests that their results are relatively suboptimal and that their robustness is weaker.
More performance details (such as the average values, standard deviations, Friedman means, etc.) are provided in the appendix.
Figure 4. Box plots of QHDE and other algorithms on CEC 2022 functions.
Six subfigures showing boxplots. QHDE generally shows smaller medians and tighter interquartile ranges.
4.3. Portfolio selection
@@
In this section, QHDE was applied to four portfolio selection problems to verify its superiority in portfolio optimization. All the stock data used in the experiments are derived from the CSI 300 Index jointly released by the Shanghai Stock Exchange and the Shenzhen Stock Exchange. The experimental data comprised the closing prices of these stocks from October 8, 2024, to November 13, 2024. We conducted a preliminary analysis of these stock data and selected the best 20, 40, 60, and 80 stocks from the 300 stocks respectively based on the average daily return during this period to form four portfolio selection problems. To eliminate bias, all eight algorithms were provided with the same computational basis, with a population size of 50 and 100 iterations. We evaluated the performance of these algorithms using three key metrics: the objective function F(E) in Equation (12), the Sharpe ratio (Sr) in Equation (8), and the equality constraint S(E) in Equation (10). The results are shown in Table 2, where the best-performing algorithm among the eight comparison algorithms is highlighted in bold.
@@
Table 2. Four portfolios’ results solved by seven metaheuristic algorithms
@@
F(E) Sr S(E)3 F(E) Sr S(E)3 F(E) Sr S(E)3 F(E) Sr S(E)3
SASS -1.99E+04 13.25 1.141 -1.65E+06 5.126 2.284 -6.72E+06 2.622 3.593 -1.26E+07 0.974 4.550
COLSHADE 15.39 15.49 1.000 -6.27E+05 2.807 1.792 -1.54E+07 2.461 4.920 -1.12E+08 0.700 11.56
sCMAES 17.63 17.63 1.000 -2.71E+06 4.424 2.647 -2.94E+07 1.355 6.426 -9.33E+07 0.704 10.66
HSEPSO 11.55 11.562 1.000 11.10 11.11 1.000 3.362 3.363 1.000 -6.18E+05 1.896 1.786
NDSOT 4.00E+06 1.699 3.000 6.40E+07 0.852 9.000 1.96E+08 0.573 15.00 4.84E+08 0.467 23.00
DE 20.58 21.04 0.999 10.11 14.718 0.998 -1.76E+06 2.974 2.328 -2.08E+07 1.515 5.555
ADE 19.94 19.96 1.000 17.66 17.80 1.000 9.703 10.50 1.001 4.464 6.644 0.999
QHDE 26.66 26.66 1.000 22.89 23.10 1.001 16.83 16.91 1.000 7.536 7.750 1.000
@@
The results presented in Table 2 indicate that, across four portfolio optimization tasks of different scales, QHDE consistently exhibits remarkable and stable comprehensive performance. The F(E) values of QHDE in the four portfolios are 26.66, 22.89, 16.83, and 7.53, representing the highest scores for each respective scale. Notably, QHDE achieves performance gains of 29.5%, 29.6%, 73.4%, and 68.8% over the runner-up, thereby demonstrating its outstanding capability in maximizing returns. Furthermore, the Sr of QHDE consistently surpasses that of other comparative algorithms, indicating its effectiveness in balancing risk and return to achieve superior risk-adjusted returns. It is worth noting that in small-scale tasks, the optimal value obtained by DE ranks second only to QHDE, suggesting that DE possesses a certain degree of exploitation capability in complex portfolio problems. However, as the scale increases, the performance of DE deteriorates rapidly. Specifically, its S(E) value at 60 dimensions is significantly higher than expected, revealing its limitations in handling complex constraints. In contrast, the proposed QHDE maintains values extremely close to 1 across all test scales. It not only improves optimization precision but also effectively resolves the deficiencies of DE in constraint handling, thereby exhibiting superior comprehensive performance.
Upon observing Table 2, it can be noted that as the scale of the stock portfolio increases, the Sr gradually decreases. This is associated with a wider range of stock selections. The overall average expected return of 80 stocks is lower than that of the optimal 20-stock portfolio. This phenomenon is consistent with real investment environments, demonstrating the reliability and robustness of the model proposed in this paper.
(a)                  (b)                 (c)
@@
Figure 5. Evaluation and comparison of QHDE applied to 20 stocks alongside other metaheuristic algorithms. (a) illustrates the overall objective function value F(E), (b) presents the Sharpe ratio (Sr), and (c) presents the equality constraint S(E).
Evaluation and comparison of QHDE applied to 80 stocks alongside other metaheuristic algorithms
To further analyze the performance of the algorithms, Figure  5 presents the iterative process of each algorithm. Constrained by the length of the paper, only the iterative process of 20 stocks is presented in the main text. The convergence details of other scales can be found in the appendix. As shown in Figure  5(a), QHDE achieves the highest F(E) value and converges the fastest, highlighting its significant advantage in global search and convergence efficiency. Other algorithms lagged significantly in both speed and optimization quality. Furthermore, DE achieves final convergence after approximately 50 iterations. Although its optimal value ranks second only to QHDE, its overall performance is markedly inferior. This reaffirms that DE’s limited convergence speed hampers its exploitation potential in complex portfolios. In contrast, by incorporating quantum computing strategies, QHDE significantly accelerates the convergence process and enhances optimization precision.
Figure  5(b) displays the Sr of the algorithms. QHDE still achieves the highest Sr, with a stable convergence process. Although its convergence speed is slightly slower than ADE, its final optimization result significantly outperforms the other algorithms. This indicates that QHDE is capable of consistently providing high returns while controlling risk, demonstrating good risk-return balance.
Figure  5(c) shows the convergence of the equality constraint S(E). QHDE exhibits a significant advantage in handling equality constraints, rapidly approaching 1 during the initial iterations, whereas other competing algorithms require dozens of iterations to approach the feasible region, with some (e.g., SASS, NDSOT) failing to converge entirely. The magnified view further confirms that QHDE maintains an error-free and stable state throughout the entire process, demonstrating its outstanding reliability in complex constraint environments.
@@
In summary, QHDE, as a hybrid optimization algorithm based on quantum evolutionary strategies, demonstrates significant advantages in complex portfolio optimization tasks. It not only significantly improves the algorithm’s global exploration and local exploitation capabilities but also exhibits strong adaptability and reliability when facing high-dimensional complex constraint problems. These characteristics make QHDE an effective tool for solving complex financial optimization problems and providing an efficient and precise solution to portfolio selection problems.
5. Conclusion

### v3

@@ 3.2.2 Potential-driven dynamic quantum tunneling strategy
3.2.3 Dynamic elite pool with Cauchy-Gaussian mixed disturbance strategy
@@
In later iterations, DE populations often cluster around the current best solution. This rapid assimilation causes stagnation and premature convergence if that solution is a local optimum. To counter this, we propose a dynamic elite pool ($Elite_{p}$) combined with a Cauchy-Gaussian mixed disturbance strategy. By maintaining a pool of top-performing individuals and applying stochastic perturbations, the algorithm preserves population diversity and strengthens its global search capacity. The specific procedure is as follows:
Figure 1: Flow chart of QHDE
@@
First, the three highest fitness individuals, denoted as $PfZ1$, $PZ2$, and $PZ3$, are selected from the population. Then the arithmetic mean of their positions is calculated to obtain the average position $PZm$:
$PZm=\frac{(PZ1+PZ2+PZ3)}{3}$ (17)
@@
Cauchy-Gaussian mutation is a hybrid mutation method that combines the long-tail characteristics of the Cauchy distribution with the concentration tendency of the Gaussian distribution. The purpose of this strategy is to apply strategic random perturbations to the global best individual, exploring potentially unknown beneficial regions around the global optimum, while maintaining the "memory" of the current best solution. After combining the dynamic elite pool and Cauchy-Gaussian perturbation strategies, the update formula for the new individual’s position is as follows:
$x^{t+1}_{\text{best}}=Elite\times[\rho\times\text{Cauchy}(0,1)+(1-\rho)\times\text{Gauss}(0,1)]$ (18)
where $x^{t+1}_{\text{best}}$ represents the global best individual in $t+1$ generation, and $\rho\in[0,1]$ denotes a weight coefficient balancing the Cauchy and Gaussian distributions.
@@
7:   Randomly select $x_{r_{1}},x_{r_{2}},x_{r_{3}}$ from the population and calculate mutant vector $v_{i}$ using (13)
@@
9:   Calculate the tunneling probability $P_{t}(x_{i})$ using (21)
10:   Generate $x_{new}$ using (22)
11:   if $x_{new}$ surpasses $u_{i}$ then
12:     $u_{i}=x_{new}$
13:   end if
14:   if $f(u_{i})\leq f(x_{i})$ then
15:     Insert $u_{i}$ into new generation $P_{new}$
16:   else
17:     Insert $x_{i}$ into new generation $P_{new}$
18:   end if
19:  end for
@@
20:  Select the fittest individuals $PZ1,PZ2,PZ3$ and calculate PZmPZm using (23) to form the elite pool
21:  Generate $x_{\text{best}}^{t+1}$ using (24)
22:  if $x_{\text{best}}^{t+1}$ surpasses $pos_{\text{best}}$ then
23:   $pos_{\text{best}}=x_{\text{best}}^{t+1}$
24:  end if
25:  Update best global position $pos_{\text{best}}$ and its fitness value $fit_{\text{best}}$
26:  $P=P_{new}$
27:  $t=t+1$
28:end while
29:return $pos_{\text{best}}$ and $fit_{\text{best}}$
3.3 Time and space complexity analysis
@@
The computational complexity of QHDE mainly depends on population initialization and core operations such as fitness evaluation and iterative updates. With population size $N$, iteration limit $T_{\max}$, and dimension $D$, the original complexity is $O(T_{\max}\times N\times D)$. Initialization operations (good-point-set generation, chaotic mapping, and reverse learning) each cost $O(N\times D)$ and are executed once. Quantum tunneling updates and the elite-pool disturbance strategy both require $O(N\times D)$ per iteration, giving $O(T_{\max}\times N\times D)$. Thus, the total complexity remains $O(T_{\max}\times N\times D)$, consistent with the original algorithm.
@@
The space complexity of DE depends on storing the population, fitness values, and candidate solutions. With population size $N$ and dimension $D$, storing positions and temporary candidates requires $O(N\times D)$, while fitness values need $O(N)$. Thus, the overall complexity is $O(N\times D)$. In the improved QHDE, storing good-point sets, reverse individuals, and the elite pool also requires $O(N\times D)$, so the total space complexity remains $O(N\times D)$.
4 Experiments and comprehensive analysis
@@
This section provides a detailed discussion of the experimental results obtained by applying QHDE to numerical optimization tasks and portfolio selection problems. We first use the CEC 2020 [8] and CEC 2022 [11] suites to test its efficacy on multimodal and high-dimensional functions. Subsequently, QHDE is applied to four portfolio problems (20, 40, 60, and 80 stocks) and compared against seven widely-used metaheuristic algorithms to validate its superiority. The simulations were conducted on a Windows 11 platform with a 64-bit operating system. The analyses were carried out using MATLAB 2023b on a system powered by an AMD Ryzen 7 4800H CPU at 2.30 GHz and 16 GB of RAM.
4.1 Test functions and comparison algorithms
@@
To evaluate QHDE, we utilized the CEC 2020 and 2022 test suites. These benchmarks encompass diverse, high-dimensional, and multimodal optimization problems that effectively mirror the non-convex objectives and complex constraints of real-world portfolio selection. Their emphasis on balancing multiple local optima and risk-return trade-offs provides a rigorous, representative basis for assessing the algorithm’s effectiveness and robustness in complex environments.
@@
The performance of QHDE was systematically compared against seven well-known algorithms. These are categorized into two distinct groups for a structured analysis: (1) State-of-the-art algorithm: SASS [7], COLSHADE [5], sCMAES [12], HSEPSO [16], NDSOT [4]; (2) DE [3] and its improved variants ADE [1]. Table 1 provides an overview of the parameter settings for eight algorithms.
Table 1: Parameter configurations for competing algorithms
@@
QHDE $\beta$, $P_{c}$, $V_{0}$, $\gamma$ 0.5, 0.1, 0.5, 0.1
4.2 Quantitative evaluation
@@
This section evaluates QHDE’s performance using the CEC 2022 test suite (CEC 2020 results detailed in the Appendix [18]). To ensure statistical reliability, we standardized the experimental parameters: a population size of 30, a maximum of 500 iterations, and 30 independent runs for each algorithm.
@@
Figure 2 presents the Friedman rank radar plot of all participating algorithms. As evident in Figure 2(a), QHDE (red) consistently occupies the central region, indicating superior (lower) rankings. As dimensionality increases to 20 in Figure 2(b), although the QHDE curve shows a broader distribution due to heightened problem complexity, it remains significantly more compact than its competitors. Notably, QHDE ranks first on functions F2, F3 and F7-12, which are known for their high complexity and multimodality. The overall rank of QHDE is the lowest among all contenders, further confirming its superior and robust optimization capability across a wide range of scenarios.
(a)                          (b)
Figure 2: Ranking distribution of different algorithms on CEC 2022. (a) Dim = 10, (b) Dim = 20.
@@
As shown in Figure 3, the boxplot analysis of various algorithms on representative functions reveals a significant stability advantage of QHDE. The box of the QHDE boxplot is relatively narrow, indicating a highly concentrated data distribution. This suggests that, across multiple experimental runs, the results generated by QHDE exhibit minimal dispersion, and the search process remains stable. By leveraging its three integrated mechanisms, QHDE effectively avoids local optima to provide stable, high-quality solutions. In contrast, for algorithms such as SASS, HSEPSO and NDSOT, the overall positions of the boxplots are relatively higher. This suggests that their results are relatively suboptimal and that their robustness is weaker.
Figure 3: Box plots of QHDE and other algorithms on CEC 2022 functions.
4.3 Ablation experiments
QHDE incorporates three enhancement strategies: the good point set-chaos reverse learning initialization mechanism (Strategy 1), the potential-driven dynamic quantum tunneling strategy (Strategy 2), and the dynamic elite pool combined with Cauchy-Gaussian mixed perturbation (Strategy 3). To examine the individual and combined effects of these strategies on QHDE’Ds performance, ablation experiments were conducted. We designed six variants (QHDE1, QHDE2, QHDE3, QHDE12, QHDE13, and QHDE23) to isolate these components. Figure 4 display the Friedman rankings on CEC 2020 (Dim=10, 20), among which the improvement of QHDE is the most remarkable.
Figure 4: Ablation experiment on CEC 2020.
Figure 4 demonstrate that QHDE achieves the highest average Friedman ranking, consistently placing in the top 5 across all test functions. Among the three strategies, Strategy 2 (quantum tunneling) serves as the primary performance driver, as variants incorporating it (QHDE2, 12, 23) consistently outrank others. While Strategy 1 (initialization) offers limited independent gains, Strategy 3 (mixed perturbation) shows increased efficacy in high-dimensional scenarios. Its synergy with Strategy 2 enables QHDE and QHDE23 to maintain superior optimization stability and global search capability as problem complexity scales.
4.4 Portfolio selection
@@
This section evaluates QHDE across four portfolio selection problems (comprising 20, 40, 60, and 80 stocks) to verify its optimization superiority. The dataset includes closing prices from the CSI 300 Index (Shanghai and Shenzhen Stock Exchanges) between January 1, 2024, and December 31, 2024. Stocks were selected based on their average daily returns to form the test portfolios. To ensure a fair comparison, all eight algorithms utilized a population size of 50 over 100 iterations. Performance was measured using three key metrics: the objective function F(E), the Sharpe ratio Sr, and the equality constraint S(E). The results are shown in Table 2, with the best performers highlighted in bold.
@@
Table 2: Four portfolios’ results solved by seven metaheuristic algorithms
@@
F(E) Sr S(E) F(E) Sr S(E) F(E) Sr S(E) F(E) Sr S(E)
SASS -2.37E+04 4.26E+00 1.15E+00 -6.49E+05 3.49E+00 1.81E+00 -5.36E+06 2.42E+00 3.32E+00 -1.43E+07 1.58E+00 4.79E+00
COLSHADE 1.59E+01 1.60E+01 1.00E+00 -5.27E+05 6.40E+00 1.73E+00 -1.59E+07 3.19E+00 4.99E+00 -1.57E+08 6.14E-01 1.35E+01
sCMAES 1.93E+01 2.02E+01 1.00E+00 -6.66E+06 3.10E+00 3.58E+00 -1.67E+07 1.66E+00 5.08E+00 -8.43E+07 8.17E-01 1.02E+01
HSEPSO 1.25E+01 1.25E+01 10.00E-01 1.12E+01 1.12E+01 1.00E+00 -1.67E+05 5.57E+00 1.41E+00 -1.80E+06 3.00E+00 2.34E+00
NDSOT 9.00E+00 9.00E+00 1.00E+00 -1.60E+07 7.14E-01 5.00E+00 -2.89E+08 4.89E-01 1.80E+01 -5.29E+08 4.42E-01 2.40E+01
DE 1.36E+01 1.45E+01 9.99E-01 6.87E+00 7.09E+00 1.00E+00 -1.51E+06 3.62E+00 2.23E+00 -2.14E+07 1.29E+00 5.63E+00
ADE 8.19E+00 1.04E+01 9.98E-01 7.74E+00 7.88E+00 1.00E+00 8.53E+00 8.53E+00 1.00E+00 3.03E+00 3.07E+00 10.00E-01
QHDE 2.81E+01 2.87E+01 1.00E+00 1.63E+01 1.71E+01 1.00E+00 1.23E+01 1.37E+01 9.99E-01 5.96E+00 1.02E+01 1.00E+00
@@
Table 2 demonstrates that QHDE maintains remarkable and stable performance across all portfolio scales. Its objective values ($F(E)$) of 28.1, 16.3, 12.3, and 5.96 are the highest in their respective categories, outperforming the runners-up by margins of 45.5%, 45.5%, 44.1%, and 96.6%. Furthermore, the Sr of QHDE consistently surpasses that of other comparative algorithms, indicating its effectiveness in balancing risk and return to achieve superior risk-adjusted returns. It is worth noting that in small-scale tasks, the optimal value of DE ranks in the second tier, suggesting that DE possesses a certain degree of exploitation capability in complex portfolio problems. However, as the scale increases, the performance of DE deteriorates rapidly. Specifically, its S(E) value at 60 dimensions is significantly higher than expected, revealing its limitations in handling complex constraints. In contrast, the proposed QHDE maintains values extremely close to 1 across all test scales. It not only improves optimization precision but also effectively resolves the deficiencies of DE in constraint handling, thereby exhibiting superior comprehensive performance.
@@
Figure 5: Evaluation and comparison of QHDE applied to 80 stocks alongside other metaheuristic algorithms. (a) illustrates the objective function value F(E), (b) presents the Sharpe ratio Sr, and (c) presents the equality constraint S(E).
Figure 5 presents the iterative process for 80 stocks (see Appendix [18] for other scales). As shown in Figure 5(a), QHDE achieved the highest F(E) value and had the fastest convergence speed, highlighting its significant advantages in global search and convergence efficiency. However, other algorithms like DE failed to converge to the optimal solution after 100 iterations, further demonstrating their limitations in high-dimensional optimization problems. In Figure 5(b), although ADE converges slightly faster initially, QHDE ultimately yields a superior Sr. This indicates that QHDE is capable of consistently providing high returns while controlling risk, demonstrating good risk-return balance. Crucially, Figure 5(c) highlights QHDE’s reliability in constraint handling. It rapidly approaching 1 during the initial iterations, competitors like sCMAES and NDSOT struggle or fail to converge. The magnified view confirms that QHDE maintains an error-free, stable state throughout the optimization process.
@@
In summary, QHDE demonstrates a significant advantage in complex portfolio optimization. It substantially enhances both global exploration and local exploitation, while maintaining exceptional adaptability and reliability when addressing high-dimensional, constrained problems. These characteristics establish QHDE as a robust and precise tool for solving complex financial optimization tasks and portfolio selection problems.
5 Conclusion